OpenStudy (anonymous):

cos φ = 4/5 with 0° < φ < 90° (a) sin 2φ

OpenStudy (anonymous):

how do you solve this??

OpenStudy (anonymous):

i know the answe is 24/25

OpenStudy (anonymous):

how do you know the answer is 24/25? and what's with the a) (sin 2phi) ?

OpenStudy (anonymous):

the a has no part in it

OpenStudy (anonymous):

okay so do you want me to solve this algebraically or tell you how to solve it with a calulator?

OpenStudy (anonymous):

actually i don't know if i know how to do with without a calculator lol

OpenStudy (anonymous):

sin2phi=2sinphicosphi we are given cosphi=4/5 then sinphi=3/5 sin2phi=2*4/5*3/5=24/25

OpenStudy (anonymous):

cos φ = 4/5 with 0° < φ < 90° (a) sin 2φ lemme just use x instead of phi... sin2x = 2sinxcosx by the double angle formula.. Look at this triangle...  this triangle is made according to the given.. So sin2x=2*(3/5)*(4/5)=24/25